OpenStudy (adamconner):

Check my work please: In the creation of aspirin, 0.998 mL of acetic anhydride is reacted with 5.166 g of salicylic acid. How many grams of aspirin do you expect to be produced from this reaction?

OpenStudy (adamconner):

so this is what I have: mass=Density* V .998x1.08=1.0778 g Acetic anhydride 1.0778/102=.01056 mol A.A 5.166g S.A/138.118= .0374 mol S.A Acetic anhydride is the L.R. so .01056x(1/1)x180.157= 1.9024 is this correct?

OpenStudy (anonymous):

looks right..

OpenStudy (adamconner):

@aaronq is this correct. when i wrote that answer it said its wrong and the correct answer is 6.74.. which means that the salicylic acid is the L.R. so are they wrong

OpenStudy (anonymous):

why don't you put the balanced reaction up for us to see?